19. (a) In an article about the cost of health care, Money magazine reported that a visit to a hospital emergency room for something as simple as a sore throat has a mean cost of $328 (Money, January 2009). Assume that the cost for this type of hospital emergency room visit is normally distributed with a standard deviation of $92. Answer the following questions about the cost of a hospital emergency room visit for this medical service.
a. What is the probability that the cost will be more than $500?
b. What is the probability that the cost will be less than $250?
c. What is the probability that the cost will be between $300 and $400?
d. If the cost to a patient is in the lower 8% of charges for this medical service, what was the cost of this patient’s emergency room visit?
1
Expert's answer
2019-10-25T11:32:22-0400

Let "X=" the cost for this type of hospital emergency room visit. Given that


"X\sim N(\mu, \sigma^2), \mu=\$328, \sigma=\$92"

Then


"Z={X- \mu\over \sigma}\sim N(0, 1)"

a. What is the probability that the cost will be more than $500? 


"P(X>500)=1-P(X\leq500)="

"=1-P(Z<{500- 328\over 92})\approx1-0.9692=0.0308"

b. What is the probability that the cost will be less than $250? 


"P(X<250)=P(Z<{250- 328\over 92})\approx0.1983"

c. What is the probability that the cost will be between $300 and $400? 


"P(300<X<400)=P(Z<{400- 328\over 92})-P(Z<{300- 328\over 92})\approx"

"\approx0.783072-0.380431=0.402641"

d. If the cost to a patient is in the lower 8% of charges for this medical service, what was the cost of this patient’s emergency room visit?


"p=0.08"

"P(Z\leq z)=0.08=>z\approx-1.405071"

"{X- 328\over 92}=-1.405071"

"X\approx\$ 198.73"